# Self-Modeling Robots by Photographing

### v1

@@ 3.2 Static reconstruction and neural bone
Inspired by BANMo (Yang et al., 2022) and BAGS (Zhang et al., 2024), we use a set of neural bones to model the kinematic structure of the robot and control the transformation of the self-model. Each bone is represented by an ellipsoid, which is defined by a center point $p\in\mathbb{R}^{3}$, a radii scaling matrix $r\in\mathbb{R}^{3}$, and a rotation expressed by quaternion $q_{b}\in\mathbb{R}^{4}$. We limit the max radii of each part with a hyperparameter. To initialize these neural bones, we use the K-means algorithm provided by Scikit-learn (Pedregosa et al., 2011) to cluster 3D Gaussians representing the static self-model. We use cluster centers to initialize center points $p_{b}$. The rotation of each bone is initialized to $(0,0,0,1)$, and the radii are initialized to half of their max value. The number of bones can be either a custom value or one plus the number of joints of the robot. All parameters of neural ellipsoid bones are automatically optimized when training. We use a kinematic network to map joint angles into bone-specific transformation, which will be explained below. In the experiment, we found that assuming all neural bones have similar volumes can get the best result. In practice, we use eight bones for the 7-DOF robot in a simulation environment and seven for the 4-DOF robot in the real world. The 4-DOF robot we use has a large base, so we use multiple bones to represent it.
@@ 3.3 Kinematic network
The kinematic network predicts how bones transform from the rest pose (canonical space) to a joint configuration pose (observation space). The structure of our kinematic network is shown in Figure. 2. Assuming there are $M$ bones, the input joints are first encoded to bone-specific code $z_{m}$ with an encoder and then decoded to rotation and translation with MLP decoder $\Psi$. We follow the idea of Tertikas et al. (2023) to build our encoder. The joint angle vector $\theta$ is first encoded with an MLP encoder $f$ and then decomposed to per-bone codes with a multi-head attention transformer $\tau$. The whole encoding process is as follows:
$z_{m}=\tau(f(\theta)),m\in{1,2,\ldots,M}.$ (4)
Then the rotation $\Delta R_{m}^{\theta}$ and translation $\Delta t_{m}^{\theta}$ of bone $m$ at joint configuration $\theta$ are decoded from the bone-specific code by the decoder $\Psi$. Using $R_{m}^{c}$ to represent the rotation matrix derived from the quaternion of bone $m$ at rest pose and using $p_{m}^{c}$ to represent the center point of bone $m$ at rest pose, the rotation matrix $R_{m}^{\theta}$ and position $p_{m}^{\theta}$ at joint configuration $\theta$ are given by:
@@
$p_{m}^{\theta}=\Delta t_{m}^{\theta}+p_{m}^{c}.$ (6)
@@
We use linear blender skinning (LBS) (Sumner et al., 2007) to calculate the transformation of each 3D Gaussian based on the bone transformation that we get from the kinematic network. We use $w_{im}^{\rightarrow}$ to represent the skinning weight of 3D Gaussian $i$ with respect to bone $m$ when transforming from rest pose to joint configuration pose. Then, the center point $\mu_{i}^{\theta}$ and the rotation matrix $R_{i}^{\theta}$ derived from the quaternion of the deformed 3D Gaussian $i$ at joint configuration $\theta$ can be computed by:
$\mu_{i}^{\theta}=\sum_{m\in M}{w_{im}^{\rightarrow}(\Delta R_{m}^{\theta}(\mu_{i}^{c}-p_{m}^{c})+p_{m}^{c}+\Delta t_{m}^{\theta})}$ (7)
@@
where $\mu_{i}^{c}$ is the center point of 3D Gaussian $i$ at the rest pose, and $R_{i}^{c}$ is the rotation matrix of 3D Gaussian $i$ at rest pose. To simplify the expression of the transformation, we use $\mu_{i}^{\theta}=\Theta^{\theta,\rightarrow}(\mu)$ to represent (7). Similarly, we also define an inverse transformation $\mu=\Theta^{\theta,\leftarrow}(\mu_{i}^{\theta})$ that is calculated as follows:
$\mu_{i}^{c}=\sum_{m\in M}{w_{im}^{\leftarrow}((\Delta R_{m}^{\theta})^{-1}(\mu_{i}^{\theta}-p_{m}^{\theta})+p_{m}^{\theta}-\Delta t_{m}^{\theta})}.$ (9)
@@
$w_{im}^{o}=(\mu_{i}-p_{m})^{T}(R_{m})^{T}r_{i}R_{m}(\mu_{i}-p_{m})$ (10)
where $r_{i}$ is a $3\times 3$ diagonal radius matrix of the 3D Gaussian $i$. When computing the weights for the transformation from the rest pose to the joint configuration pose, the distance is computed with the center points and rotation matrices of 3D Gaussian $i$ and bone $m$ at rest pose. The weights for the inverse transformation are computed with the same values at the joint configuration pose.
Similar to the approach proposed by Yang et al. (2022), delta skinning weights are used for fine geometry. Delta skinning weight $w_{im}^{\Delta}$ is given by a weight MLP $\omega$:
@@
where $\theta$ is the joint angle vector. To calculate the forward transformation weights, $\theta$ is the rest pose joint angle vector. To calculate the inverse transformation weights, $\theta$ denote the joint configuration vector.
@@
$w_{im}=softmax(-s(w_{im}^{o}+w_{im}^{\Delta}))$ (12)
where $s$ is a learnable scaling factor.
With the 3D Gaussian deformation method, we can control the static self-model. Combining the static self-model and the deformation method, we build a dynamic and controllable self-model of a robot and obtain its morphology, kinematics, and texture.
@@ 3.5 Optimization
We observed that training 3D Gaussians, the kinematic network, and neural bones simultaneously often led to convergence issues. Additionally, due to the robot’s high degrees of freedom, the kinematic network struggled to train effectively with random joint configurations. To address these challenges, after training the static 3DGS self-model, we freeze the parameters of all 3D Gaussians and focus solely on training the kinematic network. During the initial phase of training, we restrict the robot’s joint motions to a limited range by using the small-range dataset and gradually relax these constraints over time. Once the range of motion reaches its maximum, we resume joint training of the 3D Gaussian parameters and the kinematic network. In total, we train the self-model over 600,000 steps.
@@ 3.6 Loss functions
Our optimization objective $L$ is the sum of several loss functions. We use the same image reconstruction loss $L_{render}$ in 3DGS, which is a combination of $L_{1}$ and D-SSIM to compare the input with the rendering image.
@@
We also use a mask loss $L_{mask}$ that compares the foreground mask of the input image and the rendering foreground mask. $L_{mask}$ is computed by $L_{1}$ loss. Following previous work (Prokudin et al., 2023), we use the as-isometric-as-possible constrain (Kilian et al., 2007) to restrict the distance between neighboring 3D Gaussians to keep similar before and after deformation:
$L_{isopos}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}\left|d(\mu_{i}^{c}-\mu_{k}^{c})-d(\mu_{i}^{\theta}-\mu_{k}^{\theta})\right|\ }$ (14)
where $N$ is the total number of 3D Gaussians, $\mathcal{N}$ is the k-nearest neighborhood of 3D Gaussian $i$ We also use the local-rigid loss $L_{rigid}$ and the rotational loss $L_{rot}$ that is mentioned by Luiten et al. (2024) to constrain the kinematic network. Each of these losses is assigned a weight determined by the distance between 3D Gaussians:
$w_{i,k}=exp(-\lambda_{w}\|\mu_{k}^{c}-\mu_{i}^{c}\|_{2}^{2}).$ (15)
@@
$L_{rigid}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}w_{i,k}\|(\mu_{k}^{\theta}-\mu_{i}^{\theta})-R_{i}^{\theta}{(R_{i}^{c})}^{-1}(\mu_{k}^{c}-\mu_{i}^{c})\|_{2}}$ (16)
where $K$ is the number of k-nearest neighbors. The rotational loss is similarly given by:
$L_{rot}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}w_{i,k}\|R_{k}^{\theta}{(R_{k}^{c})}^{-1}-R_{i}^{\theta}{(R_{i}^{c})}^{-1}\|_{2}}.$ (17)
We use a 3D cycle loss as mentioned by Yang et al. (2022) to further constrain the kinematic network:
$L_{cycle}=\sum_{i=1}^{N}{\|\Theta^{\theta,\leftarrow}(\Theta^{\theta,\rightarrow}(\mu_{i}^{c}))-\mu_{i}^{c}\|_{2}^{2}}.$ (18)
We use a bone mask loss $L_{bone}$, a bone center loss $L_{center}$ , and a bone volume loss $L_{volume}$ to train neural bones. We treat each neural bone as a 3D Gaussian and render them into a binary image. The bone mask loss is given by the $L_{1}$ loss of the bone binary image and the robot’s mask. To discourage neural bones from being concentrated in a small region of the robot, we penalize small distances between their centers, as proposed by Noguchi et al. (2022):
$L_{center}=\frac{1}{M^{2}}\sum_{m\neq n}{\exp(\frac{\left|p_{m}-p_{n}\right|_{2}^{2}}{2\sigma^{2}})}$ (19)
@@
$L_{volume}=\frac{1}{M(M-1)}\sum_{i=1}^{M}{\sum_{k=1}^{i}{\left|\mathcal{V}(r_{i})-\mathcal{V}(r_{k})\right|}}$ (20)
where $\mathcal{V}(\cdot)$ means the volume of the neural ellipsoid bone calculated by the radius scaling matrices.
The final loss is a weighted sum of all losses mentioned above:
$L=$ $\lambda_{1}L_{render}+\lambda_{2}L_{mask}+\lambda_{3}L_{isopos}$ (21)
$+\lambda_{4}L_{rigid}+\lambda_{5}L_{rot}+\lambda_{6}L_{cycle}$
$+\lambda_{7}L_{bone}+\lambda_{8}L_{center}+\lambda_{9}L_{volume}$
where $\lambda_{i}$ (with $i=1,2,\cdots,9$) is the weight coefficient.
@@ 4 Results
We validated the proposed self-modeling method in a simulation environment using a 7-degrees-of-freedom (7-DOF) Franka robot and in an experimental setting with a 4-DOF OpenManipulator robot. Subsequently, we demonstrated the potential of our method in downstream tasks through several examples of motion planning and inverse kinematics.
@@ 4.1 Simulation
To train the model, we followed the steps below to generate $400\times 400$ resolution images of the 7-DOF Franka robot in the Pybullet simulator (Coumans and Bai, 2016–2021). First, all joints of the robot were set to 0 radians, i.e., $\theta=[0,0,0,0,0,0,0]^{T}$ rad. A total of 100 images of the stationary robot were captured from various random view angles. Next, we actuated the robot by selecting different joint angles. Initially, we constrained all joint angles within $[-\frac{\pi}{6},\frac{\pi}{6}]$ and then expanded them to $[-\frac{\pi}{3},\frac{\pi}{3}]$. Finally, we limited the joint movement to the range of $[-\frac{\pi}{2},\frac{\pi}{2}]$. In most of our simulation, we maintained the joint angles in the final range, as this range allows the robot to perform a large number of tasks while simplifying the self-model training process. We took one image for each pose of the robot from a random view angle, and the camera parameters for each image were calculated and saved during the image generation process. A total of 10,000 images were generated: 500 images with joints constrained between $[-\frac{\pi}{6},\frac{\pi}{6}]$, another 500 images with joints constrainted between $[-\frac{\pi}{3},\frac{\pi}{3}]$, and the remaining 9,000 images with joints constrainted between $[-\frac{\pi}{2},\frac{\pi}{2}]$. Additionally, we captured 9,000 images without any joint constraints to demonstrate that our method can handle this situation effectively.
@@
The results in the simulation environment are shown in Figure. 3. Four different joint configurations were randomly generated to control the self-model, and the self-model is rendered to images from three different views, including two horizontal views and a top view. Comparing the rendered images with ground truth images, it is obvious that the self-model built with our method correctly reflects the robot’s morphology, kinematics and texture under different joint configurations.
@@
As illustrated in Figure. 4a, we assigned a distinct color to each neural bone and rendered them into an image, demonstrating the self-model’s ability to learn the robot’s kinematics. Eight neural bones were used in the self-model. They are not constrained to be linked together. The order of connection and the size of each bone are automatically determined by the self-model.
@@
Figure 4: a Bone images in different joint configurations and the corresponding rendered images. Every neural bone is assigned a color randomly and rendered to images in the same way as in 3DGS. These neural bones form the kinematic link of the robot. The bone images are identical to the robot images at each joint configuration. b Simulation results when joint ranges are not limited. The reconstruction quality is worse than the case with limited joint ranges. However, the self-model still gives reliable results on how the robot moves at different joint configurations.
The quantitative results of the above four different joint configurations are presented in Table 1. The metrics we used are peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018), which are commonly used to evaluate synthetic image quality. The quantitative results for each joint configuration were calculated using seven images, consisting of six horizontal views around the robot and one top view. We also validated the proposed method in a test set containing 2,000 images captured from random view angles around the robot, with varying joint configurations. The results of the test set indicate that our self-model can correctly synthesize images of the robot at different joint configurations and view angles. It is worth noting that larger changes in joint configurations make it more challenging to obtain an accurate self-model and the test set includes joint configurations with minor changes. Consequently, the results for the 4 example joint configurations are worse than those of the test set.
@@
Table 1: Quantitative results in simulation environment

### v2

@@ 3.2 Static reconstruction and neural bone
Inspired by BANMo (Yang et al., 2022) and BAGS (Zhang et al., 2024), we use a set of neural bones to model the kinematic structure of the robot and control the transformation of the self-model. Each bone is represented by an ellipsoid, which is defined by a center point $p\in\mathbb{R}^{3}$, a radii scaling matrix $r\in\mathbb{R}^{3}$, and a rotation expressed by quaternion $q_{b}\in\mathbb{R}^{4}$. We limit the maximum radii of each part with a hyperparameter. To initialize these neural bones, we use the K-means algorithm (Arthur and Vassilvitskii, 2007) provided by Scikit-learn (Pedregosa et al., 2011) to cluster 3D Gaussians representing the static self-model. The number of clusters is the same as the number of neural bones. We use cluster centers to initialize center points $p_{b}$. The rotation of each bone is initialized to $(0,0,0,1)$, and the radii are initialized to half of their maximum values. The number of bones can be either a custom value or one plus the number of joints of the robot. All parameters of neural ellipsoid bones are automatically optimized when training. We use a kinematic network to map joint angles into bone-specific transformation, which is explained below. In the experiment, we found that assuming that all neural bones have similar volumes can get the best result. In practice, we use eight bones for the robot with 7 degrees of freedom (DOF) in a simulation environment and seven for the 4-DOF robot in the real world. The 4-DOF robot we use has a large base, so we use multiple bones to represent it.
@@ 3.3 Kinematic network
The kinematic network predicts how the bones transform from the rest pose (canonical space) to a joint configuration pose (transformed space). The structure of our kinematic network is shown in Figure 2. Assuming there are $M$ bones, the input joints are first encoded to bone-specific code $z_{m}$ with an encoder and then decoded to rotation and translation with the MLP decoder $\Psi$. We follow the idea of Tertikas et al. (2023) to build our encoder. The joint angle vector $\theta$ is first encoded with an MLP encoder $f$ and then decomposed into per-bone codes with a multi-head attention transformer $\tau$. We want each bone to focus on different parts of the encoded joint angle vector and learn features about its movement. The whole encoding process is as follows:
$\{z_{m}\}_{m=1}^{M}=\tau(f(\theta)).$ (5)
Then the rotation $\Delta R_{m}^{\theta}$ and the translation $\Delta p_{m}^{\theta}$ of bone $m$ at joint configuration $\theta$ are decoded from the bone-specific code by the decoder $\Psi$. Using $R_{m}^{c}$ to represent the rotation matrix derived from the quaternion of bone $m$ at rest pose and using $p_{m}^{c}$ to represent the center point of bone $m$ at rest pose, the rotation matrix $R_{m}^{\theta}$ and position $p_{m}^{\theta}$ at joint configuration $\theta$ are given by
@@
$p_{m}^{\theta}=\Delta p_{m}^{\theta}+p_{m}^{c}.$ (7)
@@
We use LBS (Sumner et al., 2007) to calculate the transformation of each 3D Gaussian based on the bone transformation that we obtain from the kinematic network. We use $w_{im}^{\rightarrow}$ to represent the skinning weight of 3D Gaussian $i$ with respect to bone $m$ when transforming from the rest pose to the joint configuration pose. Then, the center point $\mu_{i}^{\theta}$ and the rotation matrix $R_{i}^{\theta}$ derived from the quaternion of the deformed 3D Gaussian $i$ at joint configuration $\theta$ can be computed by
$\mu_{i}^{\theta}=\sum_{m\in M}{w_{im}^{\rightarrow}(\Delta R_{m}^{\theta}(\mu_{i}^{c}-p_{m}^{c})+p_{m}^{c}+\Delta p_{m}^{\theta})},$ (8)
@@
where $\mu_{i}^{c}$ is the center point of 3D Gaussian $i$ at the rest pose, and $R_{i}^{c}$ is the rotation matrix of 3D Gaussian $i$ at the rest pose. To simplify the expression of the transformation, we use $\mu_{i}^{\theta}=\Theta^{\theta,\rightarrow}(\mu_{i}^{c})$ to represent (8). Similarly, we also define an inverse transformation $\mu_{i}^{c}=\Theta^{\theta,\leftarrow}(\mu_{i}^{\theta})$ that is calculated as follows:
$\mu_{i}^{c}=\sum_{m\in M}{w_{im}^{\leftarrow}\left(\left(\Delta R_{m}^{\theta}\right)^{-1}\left(\mu_{i}^{\theta}-p_{m}^{\theta}\right)+p_{m}^{\theta}-\Delta p_{m}^{\theta}\right)}.$ (10)
@@
$w_{im}^{o}=(\mu_{i}-p_{m})^{T}R_{m}^{T}r_{i}R_{m}(\mu_{i}-p_{m})$ (11)
where $r_{i}$ is a $3\times 3$ diagonal radius matrix of the 3D Gaussian $i$. When computing the weights for the transformation from the rest pose to the joint configuration pose, the distance is computed with the center points and rotation matrices of 3D Gaussian $i$ and bone $m$ at the rest pose. The weights for the inverse transformation are computed with the same values at the joint configuration pose.
Similar to the approach proposed by Yang et al. (2022), delta skinning weights are used for fine geometry. Delta skinning weight $w_{im}^{\Delta}$ is given by a weight MLP $\omega$, which takes the joint angle vector and the center position of a 3D Gaussian as input, and outputs a delta weight for each bone:
@@
where $\theta$ is the joint angle vector. To calculate the forward transformation weights, $\theta$ is the rest pose joint angle vector and can be expressed as $\theta^{c}$. To calculate the inverse transformation weights, $\theta$ denotes the joint configuration vector and can be expressed as $\theta^{t}$.
@@
$w_{im}=\text{softmax}(-s(w_{im}^{o}+w_{im}^{\Delta}))$ (13)
where $s$ is a learnable scaling factor. A demonstration of the 3D Gaussian deformation is illustrated in Figure 3.
Figure 3: Illustration of 3D Gaussian deformation. We use one 3D Gaussian and two neural bones as an example. The deformation of the 3D Gaussian is a weighted sum of the transformation of the two bones. The weight is divided into two parts. One is based on the distance between 3D Gaussian and neural bones, and the other is predicted by a weight MLP.
With the 3D Gaussian deformation method, we can control the static self-model. Combining the static self-model and the deformation method, we build a dynamic and controllable self-model of a robot and obtain its morphology, kinematics, and surface color.
@@ 3.5 Optimization
We optimize 3D Gaussians, parameters of the kinematic network, parameters of all neural bones, parameters of the weight MLP, and the scaling factor for the skinning weight. We observed that training 3D Gaussians, the kinematic network, and neural bones simultaneously often led to convergence issues. To address these challenges, after training the static 3DGS self-model, we freeze the parameters of all 3D Gaussians and focus solely on training the kinematic network. Additionally, due to the robot’s high degrees of freedom, it exhibits a wide range of postural variations with random joint configurations, and the kinematic network struggles to train effectively. Therefore, during the initial training phase, we restrict the robot’s joint motions to a limited range by using the small-range dataset and gradually relax these constraints over time. This approach enables the model to learn kinematics progressively, starting from simpler to more complex instances. Once the range of motion reaches its maximum, we resume joint training of the 3D Gaussian parameters and the kinematic network. In total, we trained the self-model in over 600,000 steps with a joint limitation of $\frac{\pi}{2}$. Another 300,000 steps are required to train the self-model without joint limitation.
@@ 3.6 Loss functions
Our optimization objective $L$ is the sum of several loss functions. Our loss functions can be separated into four parts and used to train different parts of our method, including the image part, the 3D Gaussian part, the skinning weight part, and the neural bone part. For the image part, we use the same image reconstruction loss $L_{render}$ in 3DGS, which is a combination of the $L_{1}$ loss and a D-SSIM term to compare the input with the rendered image:
@@
We also use a mask loss $L_{mask}$ that compares the foreground mask of the input image and the rendered foreground mask, which is calculated as the following $L_{1}$ loss:
[…]
where $I_{m}$ is the ground truth mask and $\hat{I}_{m}$ is the rendered mask.
Then, for the 3D Gaussian part, following the previous work (Prokudin et al., 2023), we use the as-isometric-as-possible constrain (Kilian et al., 2007) to restrict the distance between neighboring 3D Gaussians to keep similar before and after deformation:
$L_{isopos}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}\frac{\left|d(\mu_{i}^{c}-\mu_{k}^{c})-d(\mu_{i}^{\theta}-\mu_{k}^{\theta})\right|}{KN}}$ (16)
where $N$ is the total number of 3D Gaussians and $\mathcal{N}$ is the set of $K$ nearest neighborhoods of 3D Gaussian $i$.
We also use the local-rigid loss $L_{rigid}$ and the rotational loss $L_{rot}$ (Luiten et al., 2024) to constrain the kinematic network. Each of these losses is assigned a weight determined by the distance between 3D Gaussians:
$w_{i,k}=\exp(-\lambda_{w}\|\mu_{k}^{c}-\mu_{i}^{c}\|_{2}^{2}).$ (17)
@@
$L_{rigid}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}\frac{w_{i,k}\|(\mu_{k}^{\theta}-\mu_{i}^{\theta})-R_{i}^{\theta}{(R_{i}^{c})}^{-1}(\mu_{k}^{c}-\mu_{i}^{c})\|_{2}}{KN}}.$ (18)
This loss forces all nearby Gaussians of Gaussian $i$ to move in the same way as it. Similarly, the rotational loss is given by
$L_{rot}=\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}\frac{w_{i,k}\|R_{k}^{\theta}{(R_{k}^{c})}^{-1}-R_{i}^{\theta}{(R_{i}^{c})}^{-1}\|_{2}}{KN}}.$ (19)
This loss ensures that all nearby Gaussians of Gaussian $i$ have the same rotation at different joint configurations. One could refer to Luiten et al. (2024) for more details about these two loss functions.
For the skinning weight part, we use a 3D cycle loss (Yang et al., 2022) to encourage the delta skinning weight that is calculated by (12) under different joint configurations:
$L_{cycle}=\frac{1}{N}\sum_{i=1}^{N}{\|\Theta^{\theta,\leftarrow}(\Theta^{\theta,\rightarrow}(\mu_{i}^{c}))-\mu_{i}^{c}\|_{2}^{2}}.$ (20)
A smooth loss is also employed to encourage neighboring 3D Gaussians to have similar skinning weights:
$L_{smooth}=\frac{1}{KN}\sum_{i=1}^{N}{\sum_{k\in\mathcal{N}}}|w_{i}-w_{k}|.$ (21)
Finally, we use a bone mask loss $L_{bone}$, a bone center loss $L_{center}$, and a bone volume loss $L_{volume}$ to train neural bones. We treat each neural bone as a 3D Gaussian and render it into a binary image. The bone mask loss is given by the $L_{1}$ loss of the bone binary image and the robot’s mask. To discourage neural bones from being concentrated in a small region of the robot, inspired by Noguchi et al. (2022), we penalize small distances between their centers:
[…]
@@
$L_{volume}=\frac{2}{M(M-1)}\sum_{j=1}^{M}{\sum_{k=1}^{j-1}{\left|\mathcal{V}(r_{j})-\mathcal{V}(r_{k})\right|}}$ (23)
where $\mathcal{V}(\cdot)$ is the volume of the neural ellipsoid bone calculated by the radius scaling matrices.
The final loss is a weighted sum of all loss functions mentioned above:
$L=$ $\lambda_{1}L_{render}+\lambda_{2}L_{mask}$ (24)
$+\lambda_{3}L_{isopos}+\lambda_{4}L_{rigid}+\lambda_{5}L_{rot}$
$+\lambda_{6}L_{cycle}+\lambda_{7}L_{smooth}$
$+\lambda_{8}L_{bone}+\lambda_{9}L_{center}+\lambda_{10}L_{volume}$
where $\lambda_{i}$ (with $i=1,2,\cdots,10$) is the weight coefficient. The first, second, third, and fourth lines represent the losses for images, 3D Gaussians, skinning weights, and neural bones, respectively.
3.7 Implementation details
For the static reconstruction, we randomly generated a point cloud with 50,000 points as the base of 3D Gaussians. We set the densification interval of 3DGS to 200 to constrain the total number of 3D Gaussians during optimization.
We used an MLP with three layers of 128 dimensions for the kinematic network to encode the input joint values. We used a transformer encoder with four heads and two layers to encode the part-specific feature. The MLP decoder has eight layers with 128 dimensions. We used quaternions to represent rotation. For the weight MLP in 3D Gaussian deformation, we used an MLP with six layers of 256 dimensions.
In terms of optimization, we used the same setup with the original implementation of 3DGS (Kerbl et al., 2023) to train 3D Gaussians. We used AdamW (Loshchilov and Hutter, 2019) to optimize the rest part of our model. We set the learning rate to $1\times 10^{-4}$ and use a cosine annealing schedule (Loshchilov and Hutter, 2017). The weight coefficients for all loss functions were empirically set as follows: $\lambda_{1}=1,\lambda_{2}=0.1,\lambda_{3}=1,\lambda_{4}=0.1,\lambda_{5}=0.1,\lambda_{6}=1,\lambda_{7}=10,\lambda_{8}=0.1,\lambda_{9}=0.1,\lambda_{10}=1$. The training process of 600,000 steps took about 16 hours on a single Nvidia RTX 3090 GPU to learn a self-model with joint limitation, and the additional 300,000 steps required to train a self-model without joint limitation took about 8 hours. Besides, we replaced the 3DGS backbone in our method with NeRF, and the NeRF version method required approximately two days of training on a single RTX 3090 GPU, yet produced significantly inferior rendering quality compared to our method.
@@ 4 Results
We validated the proposed self-modeling method in a simulation environment using a 7-DOF Franka robot and in an experimental setting with a 4-DOF OpenManipulator robot. Subsequently, we demonstrated the potential of our method in downstream tasks through several examples of motion planning and inverse kinematics. All of our experiments are conducted on an Nvidia RTX 3090 GPU, and the per-image rendering time is approximately 0.08 seconds.
@@ 4.1 Simulation
To train the model, we followed the steps below to generate $400\times 400$ resolution images of the 7-DOF Franka robot in the Pybullet simulator (Coumans and Bai, 2016–2021). First, all joints of the robot were set to 0 radians, i.e., $\theta=[0,0,0,0,0,0,0]^{T}$ rad. A total of 100 images of the stationary robot were captured from various random view angles. Next, we actuated the robot by selecting different joint angles. Initially, we constrained all joint angles within $\frac{\pi}{6},\frac{\pi}{6}]$ and then expanded them to $\frac{\pi}{3},\frac{\pi}{3}]$. Finally, we limited joint movement to the range of $\frac{\pi}{2},\frac{\pi}{2}]$. For each range, we randomly sampled seven joint angles within it and moved the simulated robot to the sampled pose. When the robot was stable, we recorded the actual joint angles attained, in case some randomly generated target configurations might be kinematically unreachable, as some configurations would cause self-collision. In most of our simulations, we maintained the joint angles in the final range, as this range allows the robot to perform a large number of tasks and leads to better image rendering quality while simplifying the self-model training process. We took one image for each robot pose from a random view angle, and the camera parameters for each image were calculated and saved during the image generation process. A total of 10,000 images were generated: 500 images with joints constrained in $\frac{\pi}{6},\frac{\pi}{6}]$, another 500 images with joints constrained in $\frac{\pi}{3},\frac{\pi}{3}]$, and the remaining 9,000 images with joints constrained in $\frac{\pi}{2},\frac{\pi}{2}]$. Additionally, we captured 9,000 images without joint constraints to demonstrate that our method can handle this situation effectively.
@@
The results in the simulation environment are shown in Figure 4. Four different joint configurations were randomly generated to control the self-model, and the self-model is rendered into images from three different views, including two horizontal views and a top view. Comparing the rendered images with ground truth images, it is obvious that the self-model built with our method correctly reflects the robot’s morphology, kinematics and surface color under different joint configurations. Besides the rendered images, we treated centers of 3D Gaussians as point clouds, and extracted a mesh with alpha shapes (Edelsbrunner et al., 1983) because it is the most direct method for converting a point cloud into a mesh. We set the parameter $\alpha$ to 0.022. We show meshes of the robot at the rest pose and joint configuration $\theta_{3}$ in Figure 5 from five views. As we employed the original version of the 3DGS method, which inherently lacks the capability to directly generate a mesh, the quality of the meshes extracted using our approach is subject to degradation.
Figure 5: Mesh extracted from the point could formed by 3D Gaussians. We illustrate five views of the extracted meshes. They are one top-down view along with four orthogonal views: front, back, left, and right.
@@
As illustrated in Figure 6(a), we assigned a distinct color to each neural bone and rendered them into an image, demonstrating the self-model’s ability to learn the robot’s kinematics. Eight neural bones were used in the self-model. They are not constrained to be linked together, as each bone is controlled independently and positioned according to its intended location under each joint configuration of the robot. The order of connection and the size of each bone are automatically determined by the self-model.
@@
Figure 6: a Bone images in different joint configurations and rendered images at the same configuration. Every neural bone is assigned a color randomly and rendered to images in the same way as in 3DGS. These neural bones form the kinematic link of the robot. The bone images correspond to the robot images at each joint configuration. b Simulation results when joint ranges are not limited. The reconstruction quality is worse than the case with limited joint ranges. However, the self-model still gives reliable results on how the robot moves at different joint configurations.
@@
Table 1: Quantitative results in the simulation environment with joint limitations.